OpenStudy (anonymous):

Find the sum of a finite arithmetic sequence from n = 1 to n = 15, using the expression 2n + 5 315 215 32 35

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

Oh so someone is helping you already?

OpenStudy (anonymous):

You could help too xD the more the better c: Because im just as confused..

OpenStudy (anonymous):

Confused?? Where??

OpenStudy (anonymous):

Everywhere. LOL

OpenStudy (anonymous):

IVe done this 8 times, and i still dont get it.

OpenStudy (anonymous):

You have given, n = 1 from 15 and you have got one expression too.. Start by putting n = 1 in that expression.. Can you do this?

OpenStudy (anonymous):

You have : 2n + 5 here just put n = 1 and tell me what do you get?

OpenStudy (anonymous):

7.

OpenStudy (anonymous):

And similarly, put 2..

OpenStudy (anonymous):

i have to do this 15 times? ;-;

OpenStudy (anonymous):

No, after 2 just one more time..

OpenStudy (anonymous):

Okay, so 7,9,11

OpenStudy (anonymous):

Oh wait, ;o

OpenStudy (anonymous):

I have asked to do only for 2, as you have done for 3 too, so its your mistake, now you will do for n = 15, you yourself have increased your work.. :P

OpenStudy (anonymous):

i go 315 :D Well, its easier for me.

OpenStudy (anonymous):

You have solved it??

OpenStudy (anonymous):

Yeah, it isnt that hard o-o one i found the first 3, i found the pattern. then i continued it up to 15, then i added.

OpenStudy (anonymous):

n=15*

OpenStudy (anonymous):

Hey, you have to do this with using Arithmetic Progression, not by simply adding them one by one.. :)

OpenStudy (anonymous):

Well, i did get it right xD

OpenStudy (anonymous):

Yes, it is maths, this is its beauty that you get same answer but yet different ways are there to do that, and you must do in the way the question is asking you to do as.. :)

OpenStudy (anonymous):

Yeah, thats the confusing way xD

OpenStudy (jdoe0001):

hmmm did you find it?

OpenStudy (jdoe0001):

\(\bf a_1=2(1)+5\to 9\qquad a_{15}=2(15)+5\to 35 \\ \quad \\ S_{\color{brown}{ n}}=\cfrac{{\color{brown}{ n}}}{2}a_1+a_{\color{brown}{ n}}\implies S_{\color{brown}{ 15}}=\cfrac{{\color{brown}{ 15}}}{2}a_1+a_{\color{brown}{ 15}}\)